OpenStudy (anonymous):

You roll two dice. What is the probability that the sum of the dice is even or the sum of the dice is greater than 6? Make sure you only account for dice sums that are both even and greater than 6 once. A 6X6 table of dice outcomes will help you to answer this question.

OpenStudy (anonymous):

is it 1/3?

OpenStudy (anonymous):

wouldnbt it be like 1/2 (remember im not that bright so im just sayiin )

OpenStudy (a_clan):

A 6X6 table of dice outcomes will help you to answer this question.

OpenStudy (anonymous):

26/36?- not sure try and check it.

OpenStudy (anonymous):

26/36

OpenStudy (anonymous):

i got 7/36

OpenStudy (anonymous):

can't be note 1,5 if different from 5,1

OpenStudy (anonymous):

first count all posiibilities which add up to 6 or greater. you get 6+5+...2+1.Then add all possibilities where the sum is even...note odd+odd = even & ev +ev =ev

OpenStudy (anonymous):

2/3

OpenStudy (amistre64):

even or greater then 6 2,4,6,8,10,12 or 7,8,9,10,11,12; cross out the doubles to get a set of: 2,4,6,7,8,9,10,11,12 1,1.........................................1/36 1,3; 3,1; 2,2; 2,2....................4/36 1,5; 5,1; 2,4; 4,2; 3,3; 3,3 .... 6/36 1,6; 6,1; 2,5; 5,2; 3,4; 4,3 .... 6/36 2,6; 6,2; 3,5; 5,3; 4,4; 4,4 .... 6/36 3,6; 6,3; 4,5; 5,4 .................. 4/36 4,6; 6,4; 5,5; 5,5 .................. 4/36 5,6; 6,5 ................................ 2/36 6,6, 6,6 ................................ 2/36 -------- 35 /36

OpenStudy (amistre64):

left of 1,1 ... and mighta forgoten how to count too

OpenStudy (amistre64):

might be easier to see the complement; 3 and 5 1,2; 2,1 .................. 2 1,4; 4,1; 2,3; 3,2..... 4 36/36 - 6/36 = 30/36 = 15/18 = 5/6 maybe

OpenStudy (amistre64):

\begin{array}c &1 & 2 & 3 & 4 & 5 & 6 &\\ 1&\color{red}{2}&3&\color{red}{4}&5&\color{red}{6}&\color{red}{7}\\ 2&3&\color{red}{4}&5&\color{red}{6}&\color{red}{7}&\color{red}{8}\\ 3&\color{red}{4}&5&\color{red}{6}&\color{red}{7}&\color{red}{8}&\color{red}{9}\\ 4&5&\color{red}{6}&\color{red}{7}&\color{red}{8}&\color{red}{9}&\color{red}{10}\\ 5&\color{red}{6}&\color{red}{7}&\color{red}{8}&\color{red}{9}&\color{red}{10}&\color{red}{11}\\ 6&\color{red}{7}&\color{red}{8}&\color{red}{9}&\color{red}{10}&\color{red}{11}&\color{red}{12}\\ \end{array} 6 out of 36 are not even or greater then 6 30/36 match